OpenStudy (evansanchez):

The sum of five consecutive integers is 75. What is the third number? 13 14 15 16

OpenStudy (otherworldly):

what do u think it is the equation would be x + ( x+1) + ( x+ 2) + ( x+3) + (x + 4)=75

OpenStudy (jango_in_dtown):

let the numbers be x-2, x-1, x, x+1 and x+2. then (x-2)+(x-1)+x+(x+1)+(x+2)=75i.e. 5x=75 i.e. x=?

OpenStudy (princeharryyy):

let the first integer be n next n+1, next n+2, next n+3 and last n+4 add them all. n+n+1+n+2+n+3+n+4 = 75 5n +10 = 75 5n = 65 n =13 third number = n+2 = 13+2 =15

OpenStudy (jango_in_dtown):

and see x is the 3rd number too.. so it is the solution too

OpenStudy (evansanchez):

15

OpenStudy (otherworldly):

wow three different equation that is totally not confusing

OpenStudy (princeharryyy):

yes 15 @evansanchez

OpenStudy (otherworldly):

wait isn't it 14

OpenStudy (triciaal):

the beauty of math, regardless of which equation you use the answer will be the same

OpenStudy (otherworldly):

true ^^^^^^ :)

OpenStudy (jango_in_dtown):

No-one liked the elegance of the solution I posted.:P

OpenStudy (otherworldly):

lol

OpenStudy (princeharryyy):

Appreciate your method @jango_IN_DTOWN thanks for the help and that's a cool statement @triciaal

OpenStudy (jango_in_dtown):

:P

OpenStudy (princeharryyy):

:)